VIS Meeting on 2016/2/19
2015/2/19 13:30 -
Participants:Takahashi, Hirata, Raffaele, Tanaka, Mark, Fabian, Shoda, Okutomi, Fujii, Aso, Ishizaki, Akutsu
Schedule status (Aso) (Chart)
- PR3 Installation Status (Takahashi)
- When releasing the IM, one of the flags on the TM was broken (came off).
- We decided to fix the mirror with respect to the RM. Now the suspension acts as a double-pendulum, damped by the IM.
- Yaw can be adjusted by the traverser, Pitch can be adjusted by the IM.
- Transfer functions from IM to IM were measured.
- Some strange results
- The vertical resonant frequency is higher than expected
- Length transfer function has many unknown peaks
- The frame for the clean bag is bent.
- Mirapro will come on 22nd to make changes.
- Pre-installation cleaning of the suspension
- Vacuum cleaner
Wiper-> Ask Hirose-san to send to Kamioka (Aso)
Ethanol-> Bring the bottle to the tunnel.
- Crane will not be available on 2/25 in the morning for the crane upgrade to remote control.
Plan for the next week
- Mon: F2F
- Tue: Re-cabling, LVDT check, fishing rod check, traverser check.
- Wed: Vacuum installation preparation
- Thu: Vacuum installation
- Fri: Post-vacuum installation work
iKAGRA ETM suspension
- Current status (Ohishi)
- Travel (Week of 2/22)
Takahashi:Wed - Thu (Fri: Akeno for the acceptance of IP from NIKHEF)
Shoda:Tue - Fri
Okutomi:Tue - Fri
Fujii:Tue - Fri
KAGRA F2F 2/21 - 23
- Travel (Week of 2/29)
- Okutomi: ?
Other Mechanical Issues
- #Production of additional PR winches (Akutsu)
Shipping status (Takahashi, Hirata) -> Wrapped. Will be shipped with IP in the next week.
- #OSEM assembly status (Akutsu)
- #List of what should be changed for PR3 hanging
- People should share ideas whenever they come up with missing thigns.
- Shoda made a spread sheet for keep tracking the ideas
- All the items are now done or being taken care of
- #Missing items(to be purchased soon)
#Traverser motor driver (Shoda) -> Will discuss Uraguchi-san. Need a box for stepping motor driver. -> These are for PR2. Will be ordered sometime.
- What are necessary in the next few weeks ?
- Watch dog system (Tanaka)
- Ordered parts to Hokushin
- About to order PCB to P-ban
- #Electronics diagram and review
- #Need to make Guradian script for PR3. Shoda + Fujii and Okutomi will write one after the mechanical installation is finished.
- #Digital System (Okutomi, Fujii)
- Three big categories of the tasks
- Type-A: Leader = Takahashi: Task list, team members
- Type-B (BS and SRC): Leader = Mark: Task list, team members
- Type-Bp modification: Leader = Shoda (Aso, Raffaele)
- Kick off meeting within February (Mark)
- BS prism preparation (Mark)
5mm prism drawing
- Confirmed with Hirata-san that BS can use same 5 mm prism as SR (grooves, wire clearance all OK).
- Confirmed with Tatsumi-san that not a big deal to use smaller BS prism as long as secondary position not changed.
- Submitted 5 mm prism drawing to company that did LIGO HSTS prism (except grooves), waiting on response.
- Second company that did LIGO groove cutting not interested in quoting - exploring other US companies.
- Should contact Shinko-sha again (Takahashi)
- Asked a quotation to Shinko-sha, gave the new drawing. Will ask about anealing (Takahashi)
- Can we make SR wire breakers bigger ?
Check what happens if we widen the slit by 3.75mm and make a new M5 (Hirata) -> Result
- Two concerns
- The rework may bend the wing shaped parts.
- The P.C.D. of the new screw holes may not coincide exactly with the other existing holes
- Consult with Fukushima-san again. Takahashi will contact Fukushima-san
- 型振り放電加工機 may be better
- Will test machine the dirty recoil mass. Will have another meeting with Fukushima-san (Takahashi)
Do we need to widen the slits for PRC RMs ? -> Depends on how difficult it is to rework the RM.
Wider cuts will break the horizontal balance. Need to check how much compensation mass is reduced (Hirata) -> 2mm reduction will be necessary. We can do this.
- Two concerns
- Flag design (Mark)
- Make shorter by small changes
- Widen the OSEM opening (Akutsu)
- BS dummy optics design (Fabian)
- Finalizing the design by the end of next week
IP base clamp (Fabian): Reduce the number of clamps-> Hirata-san finished 2D drawings. Trying to get quotation now.-> Still waiting for the quotation.
- Bread board spring (Takahashi): Hirata-san is getting quotations.Asked quotation with the above one all at once.
- #BS Assembly Frame Design (Aso, Hirata)
- #BS gluing jigs (Tatsumi)
- #Not enough crane height problem (Hirata, Saito)
- #BS clean bag (Aso)
Task list (Takahashi)
- Need to decide members
- #Type-A Filters Optimal Load Measurements
#Takahashi will make a plan for the optimal load check and ask ATC.-> Check two filters per month from April 2016. -> Will be presented to the ATC steering committee.
- #How to manufacture the long maraging wires ?
- One prototype was machined by Ojima
Aoshima can do the heat treatment for the long rod. Will ask them soon (Takahashi) -> Still waiting for a quotation.
- How to test it ?
- We can use the prototype assembly frame in TAMA.
- PR2 Wire breaker problem
- Prism removal from PR2 (Hirose, Mark)
- We need to send the spare mirror and the containers back to Kashiwa soon.
- Bring out the spare mirror and two containers out from the tunnel on 24th.
- Need to ask someone for loading help
- Bring back CLIO suspension too (Takahashi)
- Prism removal from PR2 (Hirose, Mark)
- No incident for VIS
- Some injuries in other groups
- #GAS blade inspection by Virgo (Takahashi)
- Shipped the blades. Will get results in 1-2 months.
2015/2/26 (Fri) 13:30 -
Next F2F 2016/4/1